OpenStudy (anonymous):

1. Together, Amy, Barbara, and Stan have less than 20 baseball cards. Amy has more cards than Barbara, and Barbara has more than Stan. Amy gives Stan 3 cards, and Barbara gives him 2 cards. Then they all have the same amount of cards. How many baseball cards did each person get? 2. A log is partly buried in the mud. One third of the log is buried and 8 m of the log are above ground. What is the total length of the log? 3. At a bank, Joseph asked for change for 7 one dollar bills. The bank teller gave her an equal amount of dimes and quarters. How many of each coin did Joseph get?

OpenStudy (anonymous):

@apoorvk

OpenStudy (apoorvk):

Okay so @wangkan07 let's start with the easiest one - #2. If one-third of the log was buried, then what part of it is above the ground?

OpenStudy (anonymous):

one?

OpenStudy (anonymous):

I dont really know

OpenStudy (anonymous):

is the answer 24?

OpenStudy (apoorvk):

Well okay let's do it this way. Let the length of the log be 'x'. Now one-third of the log is buried inside, that is (1/3)*x, or x/3 is buried So, the part that remains above = x - (x/3) And this part that's above, measures 8 meters, so: x - (x/3) = 8 Can you solve this and find 'x'?

OpenStudy (apoorvk):

No it isn't 24 really.. try this^

OpenStudy (anonymous):

4?

OpenStudy (apoorvk):

Nope... I think you are just too worked up and nervous- calm down and relax, this is pretty simple and you know can solve this :) I derived and posted an equation above - try to read it again and understand it - you should be able to solve it. :)

OpenStudy (anonymous):

12?

OpenStudy (apoorvk):

Yes correct-o!!! feeling better now? :)

OpenStudy (anonymous):

yes

OpenStudy (apoorvk):

Okay now the third one... You gave 7 one dollar bills, that is 7 dollars, or 700 cents right?

OpenStudy (anonymous):

yes

OpenStudy (apoorvk):

And he got in return an equal number of dimes and quarters, say 'x' no. of each of them. So the total amount he got back was x times 10 cents plus x times 25 cents - right?

OpenStudy (anonymous):

yes

OpenStudy (apoorvk):

And the amount he got back, was obviously equal to the amount he gave to the bank-teller, that is 7 dollars, or 700 cents. So, if I formulate an equation, what would that be?

OpenStudy (anonymous):

.25x+.10x=700

OpenStudy (apoorvk):

actually, a wee bit of a mistake, since you're already using cents, you don't need to use " 0. 10" and "0.25" Your equation would just be: 25x + 10x = 700

OpenStudy (apoorvk):

(Also, in case you wanted to keep the units in dollars, then your equation would have been" 0.10x + 0.25x = 7 - which is also equally correct)

OpenStudy (anonymous):

I got the rest. Thanks for the help

OpenStudy (apoorvk):

Great, could you solve no. 1 as well?

OpenStudy (apoorvk):

@wangkan07 ??